OpenStudy (anonymous):

what is the standard form of the expression i/2+i

OpenStudy (anonymous):

\[\frac{ i }{ 2+i }?\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{ i }{ i+2 }*\frac{ i-2 }{ i-2 }\] \[\frac{ i(i-2) }{-5 }\] \[\frac{ -1 - 2i }{-5 }\] \[\frac{ 1 + 2i }{5 }\]

OpenStudy (anonymous):

thats the answer?